Scientific Operations
 Front Page
 Scientific Ops
  » Plans
  » Pointings
  » Report
  » Requests
  » Log
  » Notes
  » Tech Info
Current Weekly Report
Select date for other reports

                CDS Weekly Report for June 25 - July 1, 2001       

1.  Operations Personnel

Operations Scientist:                   
            Dominic Zarro
Operations Engineers:
            Ron Yurow, Chris Maroney            

Visitors: None

2.  Science Operations

Active region observations concentrated on three targets: AR 9511, 9512,
& 9513 that were selected in conjunction with TRACE.

Mon June 25

Performed Full Sun and NIS/GIS spectral atlas studies, followed by
Polarity Reversal studies.

Tue June 26

Performed flare diagnostic studies (ARDIAG_2 & SIG_FT44) on AR 9511 in
conjunction with TRACE. TRACE subsequently repointed from AR 9511 to AR
9513 because of low count rates. Following Polarity Reversal study, CDS
switched to observing the active region loop complex associated with AR

Wed June 27

Continued flare diagnostic studies of AR 9511 and 9513 in conjunction
with TRACE. Supported HESSI campaign observations of AR 9512, although
activity levels remained low.

Thu June 28

Spacecraft maneuvers. No observations.

Fri June 29

CDS Filament engineering. No observations.

Sat June 30

CDS Filament engineering. No observations.

Sun July 1

Performed Explosive Events study. Supported SOHO/VLA Coronal Hole
Boundary campaign at coronal hole target in SE quadrant. 

3.  Instrument Status

On June 25, an NIS Setup Register A Failure System Error was detected.

On June 27, the NIS was reset at 22:34 UT.  Following the NIS reset, the
secondary (Science) processor was reset at 22:46 UT.  CDS doors were
then closed at 23:02 UT in preparation for S/C maneuvers.

GIS Filament Dumps were performed, from 15:00 UT - 23:45 UT on June 29.

On June 30, GIS Filament Dumps were performed from 15:30 UT - 19:40 UT.  

4.  Ground System Status

The following studies were imported to the EOF:


The following Study Duration changes were made:
FSUN 	(Raster ID: 11 / Var: 7)  | Old= 549s  / New= 555s
SIG_FT44(Raster ID: 11 / Var: 90) | Old= 1502s / New= 1515s
STRE4W	(Raster ID: 11 / Var: 96) | Old= 6145s / New= 6155s
O_SPOT_A(Raster ID: 36 / Var: 5)  | Old= 1524s / New= 1530s
O_SPOT_A(Raster ID: 193/ Var: 1)  | Old= 1610s / New= 1615s
SAS250N (Raster ID: 162/ Var: 5)  | Old= 3791s / New= 3930s.

A Software Update was performed on June 27.

From the CDS Operations Management Team in the Space Science & Technology Department at CCLRC Rutherford Appleton Laboratory
Site maintained by John Rainnie.
Last revised on Tuesday (17/Jul/2001) at 08:53.